OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 1.2, 2.4, 4.8, 9.6,...

OpenStudy (anonymous):

A. an = 9.6(1/2)n-1 B. an = 1.2(2)n-1 C. an = 9.6(2)n-1 Again, they are supposed to be: \[a _{n}\] These problems just really confuse me. Sorry I am posting so much! :/

OpenStudy (mertsj):

B

OpenStudy (anonymous):

Thanks :)

OpenStudy (mertsj):

Don't you have a formula for the nth term of a geometric sequence?

OpenStudy (anonymous):

Not that I am sure of.

OpenStudy (mertsj):

\[a _{n}=a _{1}(r)^{n-1}\]

OpenStudy (mertsj):

Have you ever seen anything like that?